OpenStudy (anonymous):

HELP! What is the equation of the line in slope-intercept form? the line perpendicular to y = 1/3x+5 through (2, 1)

OpenStudy (anonymous):

y=1/3x+1/3

OpenStudy (anonymous):

Is that an answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

A. y =1/3 x + 7 B. y = -3x + 7 C. y = 3x + 7 D. y = -1/3x + 7

OpenStudy (anonymous):

those are my answers

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Thank u!